OpenStudy (smarty):

Find the inverse of the following matrix. [ 3 1] [ 4 2 ]

OpenStudy (anonymous):

very snap way of finding the inverse of a two by two matrix

OpenStudy (anonymous):

in english, switch the element in the main diagonal, take the opposite of the other two divide by the determinant

OpenStudy (anonymous):

Shortcut method is to take the determinant first i.e 3*2 - 4*1 = 6 - 4 = 2  May be I am wrong... Please confirm it..